How do I find the third point on a line when doiing intercepts ?
My problem is
2x-8= -4y and the question asked find the intercepts and plott them.

Answer by rfer(16322)

(0,2)
(4,0)
2*2-8=-4y
4-8=-4y
-4/-4=y
y=1
(2,1)